### v2

@@ I INTRODUCTION
Here, we demonstrate a versatile and on-demand protocol for deterministic parameter estimation using a single mode with a standard bosonic cQED hardware. We show quantum-enhanced performance for phase as well as amplitude estimation using the superposition of coherent states (SCS), achieving state-of-the-art metrological gain in the low photon number regime, which minimizes the performance penalties due to hardware imperfections. We present a comprehensive analysis of key metrological figures of merit, including the dynamical range and demonstrate that our scheme can be adapted on the fly to optimize specific performance metrics. This is accomplished by varying a few programmable pulse parameters in the protocol to modify the superposition weights of the SCS. The efficiency and versatility of our implementation allow on-demand optimization of key performance metrics tailored to the metrological goal and specific hardware constraints. Finally, our strategy is adaptable across different physical systems and can be applied to other physical implementations of bosonic modes, such as optical resonators [29], trapped ions [30], neutral atoms [31], and mechanical resonators [32].
@@ IV METROLOGICAL FIGURES OF MERIT
In addition to the FI, which only sets the lower bound to the precision according to the Cramér-Rao bound, we now present a benchmark for direct calculation of the phase precision, $\Delta\theta$. It is computed as $\Delta\theta=\Delta p/|\partial p/\partial\theta|$, where $\partial p/\partial\theta$ is obtained from the polynomial fit as in Fig. 2(c) inset and $\Delta p$ is the standard deviation of the single shot measurement with $1000$ repetitions. The standard deviation for SCS (blue circles) is plotted in Fig. 3(c). The ideal precision, $\Delta\theta=1/\sqrt{4\overline{n}+4\overline{n}^{2}}$ (see Appendix C), surpasses the precision scaling $\Delta\theta=1/\overline{n}$ commonly used in optical [14, 15] or atomic [34] interferometers, as well as that achieved using superposition of Fock states ($\Delta\theta=1/2\overline{n}$[27, 28]. A linear fit of the results obtained using SCS also shows that $\Delta\theta_{\text{SCS}}\propto\bar{n}^{-0.915}$ in the range $\bar{n}=[0.07,1.76]$, which approaches the Heisenberg limit scaling $\Delta\theta\propto\bar{n}^{-1}$. To benchmark the quantum enhancement of our implementation, we compute the metrological gain of $20\log_{10}(\Delta\theta_{\text{CS}}/\Delta\theta_{\text{SCS}})=7.5(6)$ dB with a low photon number of 1.76. Here, the uncertainty is one standard deviation obtained from bootstrapping and stabilizes at $\sim 600$ measurement repetitions, see Appendix E. With only minor adjustments to the protocol, we also showcase amplitude estimation with significant quantum enhancement, see Appendix E.
@@ Appendix A Optimality of the protocol
$=$ $|\langle\psi|\psi(\beta)\rangle|^{2}$
@@
$+\beta^{4}\left(\frac{1}{12}\langle\hat{B}^{4}\rangle-\frac{1}{3}\langle\hat{B}\rangle\langle\hat{B}^{3}\rangle+\frac{1}{4}\langle\hat{B}^{2}\rangle^{2}\right)$
@@
$-\beta^{2}\left(\langle\hat{B}^{4}\rangle-4\langle\hat{B}\rangle\langle\hat{B}^{3}\rangle+3\langle\hat{B}^{2}\rangle^{2}-4(\Delta\hat{B})^{2}\right)$
@@
Figure 5: Schematic representation of 3D cQED hardware. The device comprises three essential components: the cavity functions as the bosonic mode, the transmon introduces nonlinearity, and the resonator serves for signal readout. Additionally, one pin is coupled with the cavity for driving, while the transmon and resonator share the same channel for operation.
@@
Our device consists of a standard three-dimensional (3D) cavity, an ancillary transmon and a planar readout resonator, see Fig. 5. The cavity is made of high-purity (4N) aluminum and is a $\lambda/4$ cylinder-shaped transmission line surrounded by a wall. The readout resonator is deposited together with the transmon on the sapphire using a double-angle evaporation technique. The sapphire is inserted in the tunnel, with the pad slightly extending to the cavity. The cavity serves as a high-Q quantum memory to store a bosonic mode and is designed to dispersively couple with the ancillary transmon qubit, which provides the necessary nonlinearity. The low-Q readout resonator is also dispersively coupled with the transom qubit, enabling fast single-shot measurement.
@@
Transmon anharmonicity (kHz) $\alpha/2\pi$ $175.3$
@@
2nd cavity-transmon dispersive shift (MHz) $\chi^{\prime}/2\pi$ 15.76
Transmon-resonator dispersive shift (MHz) $\chi_{qt}/2\pi$ 0.64
@@ Appendix C Preliminary investigations using different states
In the context of phase estimation, we investigate several states such as CS $|\alpha\rangle$, superposition of Fock states $(|0\rangle+|N\rangle)/\sqrt{2}$, squeezed vacuum $e^{\frac{-r}{2}(\hat{a}\hat{a}-\hat{a}^{\dagger}\hat{a}^{\dagger})}|0\rangle$, and SCS $\mathcal{N}(|0\rangle+|\alpha\rangle)$. Here, the normalisation factor $\mathcal{N}_{\alpha}=1/\sqrt{2(1+e^{-|\alpha|^{2}/2})}$. The average photon number $\overline{n}=\langle\hat{a}^{\dagger}\hat{a}\rangle$ of these states are respectively $|\alpha|^{2}$, $N/2$, $\sinh^{2}{r}$, and $\mathcal{N}_{\alpha}^{2}|\alpha|^{2}$. The QFI $4(\Delta\hat{n})^{2}$ of these states in terms of $\bar{n}$ are respectively $4\overline{n}$, $4\overline{n}^{2}$, $8\overline{n}^{2}+8\overline{n}$, and $4((1-\mathcal{N}_{\alpha}^{2})/\mathcal{N}_{\alpha}^{2})\overline{n}^{2}+4\overline{n}$. The QFI of SCS reduces to $4\overline{n}^{2}+4\overline{n}$ for large $\overline{n}$ as $\mathcal{N}_{\alpha}\approx 1/\sqrt{2}$.
@@
Figure 8: Position of maximum sensitivity. (a) Theoretical FI for SCS vs $\theta$ for states with different average photon number $\bar{n}$. The pink and red lines are FI at $\beta\rightarrow 0$ and maximum probability slope $|\partial p/\partial\theta|$ respectively. Experimental data (blue dot) and simulation (blue dash line) from Fig. 3(a) in the main text are replotted here for comparison. (b) The corresponding different FI are plotted for comparison.
@@ Appendix D Imperfections in the system
We use gradient ascent pulse engineering (GRAPE) pulse to realize the two-state transfer of the protocol. The GRAPE pulses ($\zeta(t)$ and $\epsilon(t)$) are optimized in an ideal system without imperfections. In practice, each element of the system is susceptible to environmental noise, making it crucial to account for the coherence and energy decay rate of both the cavity and transmon qubit. Table 1 presents the corresponding parameters. To account for decoherence, the simulations in the main text (all the sequence) are performed with the following master equation:
@@
where $\hat{L}_{k}$ denotes the jump operator for the corresponding decoherence. We consider five types of decoherence: qubit relaxation $\hat{L}_{1}=\hat{q}(1+E_{q})/\sqrt{T_{q,1}}$, qubit thermal excitation $\hat{L}_{2}=\hat{q}^{\dagger}E_{q}/\sqrt{T_{q,1}}$, qubit dephasing $\hat{L}_{3}=\hat{q}^{\dagger}\hat{q}/\sqrt{T_{q,2}}$, cavity relaxation $\hat{L}_{4}=\hat{a}(1+E_{c})/\sqrt{T_{c,1}}$, cavity thermal excitation $\hat{L}_{5}=\hat{a}^{\dagger}E_{c}/\sqrt{T_{c,1}}$, where $E_{c},E_{q}$ are the spontaneous excitation rate of the cavity and qubit population respectively.
@@
Figure 10: Performance of quantum-enhanced amplitude precision. (a) The average FI of SCS for which it is above the max FI of CS and $\gamma$ offset at which FI is maximum vs $n$. The dotted line is lineaer fit of $3.2(2)\overline{n}^{1.0(3)}+2.2(1)$. (b) The same protocol as in Fig. 4(c) in main text with initial states prepared with an offset. (c) The amplitude precision vs $n$ for both SCS (blue circles) and CS (purple diamonds). The dotted lines are linear fit of $\log(\Delta\gamma)=a\log(\overline{n})+b$ with $a$ and $b$ as free parameters. The fitted parameters are $a,b=-0.460(3),-1.311(2)$ for SCS and $a,b=-0.0105(2),-0.480(1)$ for CS.
@@ Appendix E Amplitude estimation using the superposition of coherent states
To illustrate the versatility of our protocol, we demonstrate quantum-enhanced precision for amplitude estimation by simply changing the process unitary of Fig. 2(a) in the main text to a displacement in the imaginary direction $\hat{U}(\gamma)=\exp(i\gamma(\hat{a}^{\dagger}+\hat{a}))$, where $\gamma$ is the displacement amplitude. The process generator is now $\hat{x}=(\hat{a}^{\dagger}+\hat{a})$, making states with high variance $(\Delta\hat{x})^{2}$ useful for amplitude estimation. Here, SCS offer high variance $(\Delta\hat{x})^{2}=2\overline{n}+1$ compared to CS, which have a constant variance $(\Delta\hat{x})^{2}=1$. The illustration is similar to that of phase estimation of Fig. 2(b) in the main text, except the SCS are now displaced in the imaginary direction instead of rotated. SCS with higher $\alpha$ have more rapidly oscillating fringes such that the overlap with the initial state reduces more sharply as $\gamma$ increases.
@@ F.2 Phase estimation
Next, we compares the uncertainty $\Delta\theta$ derived from Bayesian estimation with the Cramér-Rao bound prediction via $\Delta\theta=1/\sqrt{N\,\text{FI}}$ in Fig. 13. The close agreement between the two methods, within the dynamical range, demonstrates the effectiveness of Bayesian inference in achieving near-optimal parameter estimation. The minimal uncertainty $\Delta\theta_{\text{min}}$ as a function of the coherent state amplitude $\alpha$ (or average photon number $\bar{n}$) exhibits a decreasing trend, as shown in Fig. 14, indicating enhanced precision with larger $\bar{n}$. This scaling is consistent with theoretical predictions based on the Cramér-Rao bound.

### v3

@@ I INTRODUCTION
Here, we demonstrate a versatile and on-demand protocol for deterministic parameter estimation using a single mode with a standard bosonic cQED hardware. We show quantum-enhanced performance for phase as well as amplitude estimation using the superposition of coherent states (SCS), which minimizes the performance penalties due to hardware imperfections, achieving state-of-the-art metrological gain in the low photon number regime. We present a comprehensive analysis of key metrological figures of merit, including the dynamical range and demonstrate that our scheme can be adapted on the fly to optimize specific performance metrics. This is accomplished by varying a few programmable pulse parameters in the protocol to modify the superposition weights of the SCS. The efficiency and versatility of our implementation allow on-demand optimization of key performance metrics tailored to the metrological goal and specific hardware constraints. Finally, our strategy is adaptable across different physical systems and can be applied to other physical implementations of bosonic modes, such as optical resonators [29], trapped ions [30], neutral atoms [31], and mechanical resonators [32].
@@ IV METROLOGICAL FIGURES OF MERIT
In addition to the FI, which only sets the lower bound to the precision according to the Cramér-Rao bound, we now present a benchmark for direct calculation of the phase precision, $\Delta\theta$. It is computed as $\Delta\theta=\Delta p/|\partial p/\partial\theta|$, where $\partial p/\partial\theta$ is derived from the polynomial fit as in Fig. 2(c) inset and $\Delta p$ is the standard deviation of the single shot measurement with $1000$ repetitions. The standard deviation for SCS (blue circles) is plotted in Fig. 3(c). The ideal precision, $\Delta\theta=1/\sqrt{4\overline{n}+4\overline{n}^{2}}$ (see Appendix C), surpasses the precision scaling $\Delta\theta=1/\overline{n}$ commonly used in optical [14, 15] or atomic [34] interferometers, as well as that achieved using superposition of Fock states ($\Delta\theta=1/2\overline{n}$[27, 28]. A linear fit of the results obtained using SCS also shows that $\Delta\theta_{\text{SCS}}\propto\bar{n}^{-0.915}$ in the range $\bar{n}=[0.07,1.76]$, which approaches the Heisenberg limit scaling $\Delta\theta\propto\bar{n}^{-1}$. To benchmark the quantum enhancement of our implementation, we compute the metrological gain of $20\log_{10}(\Delta\theta_{\text{CS}}/\Delta\theta_{\text{SCS}})=7.5(6)$ dB with a low photon number of 1.76. Here, the uncertainty is one standard deviation obtained from bootstrapping and stabilizes at $\sim 600$ measurement repetitions, see Appendix E. With only minor adjustments to the protocol, we also showcase amplitude estimation with significant quantum enhancement, see Appendix E.
@@ Appendix A Optimality of the protocol
$=$ $|\langle\psi|\psi(\beta)\rangle|^{2}.$
@@
$+\beta^{4}\left(\frac{1}{12}\langle\hat{B}^{4}\rangle-\frac{1}{3}\langle\hat{B}\rangle\langle\hat{B}^{3}\rangle+\frac{1}{4}\langle\hat{B}^{2}\rangle^{2}\right).$
@@
$-\beta^{2}\left(\langle\hat{B}^{4}\rangle-4\langle\hat{B}\rangle\langle\hat{B}^{3}\rangle+3\langle\hat{B}^{2}\rangle^{2}-4(\Delta\hat{B})^{2}\right),$
@@
Figure 5: Schematic representation of a bosonic cQED hardware. The device comprises three essential components: the cavity functions as the bosonic mode, the transmon introduces nonlinearity, and the resonator serves for single-shot readout. Additionally, one pin is coupled with the cavity for driving, while the transmon and resonator share the same channel for operation.
@@
Our device consists of a standard three-dimensional (3D) cavity, an ancillary transmon and a planar readout resonator, see Fig. 5. The cavity is made of high-purity (4N) aluminum and is a $\lambda/4$ cylinder-shaped transmission line surrounded by a wall. The readout resonator is deposited together with the transmon on the sapphire using a double-angle evaporation technique. The sapphire is inserted in the tunnel, with the pad slightly extending to the cavity. The cavity serves as a high-Q quantum memory to store a bosonic mode and is designed to dispersively couple with the ancillary transmon qubit, which provides the necessary nonlinearity. The low-Q readout resonator is also dispersively coupled with the transmon qubit, enabling fast single-shot measurement.
@@
Transmon anharmonicity (MHz) $\alpha/2\pi$ $175.3$
@@
2nd cavity-transmon dispersive shift (kHz) $\chi^{\prime}/2\pi$ 15.76
Transmon-resonator dispersive shift (MHz) $\chi_{qr}/2\pi$ 0.64
@@ Appendix C Preliminary investigations using different states
In the context of phase estimation, we investigate several states such as CS $|\alpha\rangle$, superposition of Fock states $(|0\rangle+|N\rangle)/\sqrt{2}$, squeezed vacuum $e^{\frac{-r}{2}(\hat{a}\hat{a}-\hat{a}^{\dagger}\hat{a}^{\dagger})}|0\rangle$, and SCS $\mathcal{N}_{\alpha}(|0\rangle+|\alpha\rangle)$. Here, the normalisation factor $\mathcal{N}_{\alpha}=1/\sqrt{2(1+e^{-|\alpha|^{2}/2})}$. The average photon number $\overline{n}=\langle\hat{a}^{\dagger}\hat{a}\rangle$ of these states are respectively $|\alpha|^{2}$, $N/2$, $\sinh^{2}{r}$, and $\mathcal{N}_{\alpha}^{2}|\alpha|^{2}$. The QFI $4(\Delta\hat{n})^{2}$ of these states in terms of $\bar{n}$ are respectively $4\overline{n}$, $4\overline{n}^{2}$, $8\overline{n}^{2}+8\overline{n}$, and $4((1-\mathcal{N}_{\alpha}^{2})/\mathcal{N}_{\alpha}^{2})\overline{n}^{2}+4\overline{n}$. The QFI of SCS reduces to $4\overline{n}^{2}+4\overline{n}$ for large $\overline{n}$ as $\mathcal{N}_{\alpha}\approx 1/\sqrt{2}$.
@@
Figure 8: Position of maximum sensitivity. (a) Theoretical FI for SCS vs $\theta$ for states with different average photon number $\bar{n}$. The pink and red lines are FI at $\beta\rightarrow 0$ and maximum probability slope $|\partial p/\partial\theta|$ respectively. Experimental data (blue circles) and simulation (blue dash curve) from Fig. 3(a) in the main text are replotted here for comparison. (b) The corresponding different FI are plotted for comparison.
@@ Appendix D Imperfections in the system
We use gradient ascent pulse engineering (GRAPE) pulse to realize the two state-transfer operations in the protocol. The GRAPE pulses ($\zeta(t)$ and $\epsilon(t)$) are optimized in an ideal system without imperfections. In practice, each element of the system is susceptible to environmental noise, making it crucial to account for the coherence and energy decay rate of both the cavity and transmon qubit. Table 1 presents the corresponding parameters. To account for decoherence, the simulations in the main text (all the sequence) are performed with the following master equation:
@@
where $\hat{L}_{k}$ denotes the jump operator for the corresponding decoherence. We consider five types of decoherence: qubit relaxation $\hat{L}_{1}=\hat{q}\sqrt{(1+n_{q})/T_{q,1}}$, qubit thermal excitation $\hat{L}_{2}=\hat{q}^{\dagger}\sqrt{n_{q}/T_{q,1}}$, qubit dephasing $\hat{L}_{3}=\hat{q}^{\dagger}\hat{q}\sqrt{2/T_{q,\phi}}$, cavity relaxation $\hat{L}_{4}=\hat{a}\sqrt{(1+n_{c})/T_{c,1}}$, cavity thermal excitation $\hat{L}_{5}=\hat{a}^{\dagger}\sqrt{n_{c}/T_{c,1}}$, where $T_{q,\phi}=1/(1/T_{q,2}-0.5/T_{q,1})$ and $n_{c}$ ($n_{q}$) is the average thermal excitations in the cavity (qubit).
@@
Figure 10: Performance of quantum-enhanced amplitude precision. (a) The average FI of SCS for which it is above the max FI of CS and $\gamma$ offset at which FI is maximum vs $\bar{n}$. The dotted line is linear fit of $3.2(2)\overline{n}^{1.0(3)}+2.2(1)$. (b) The same protocol as in Fig. 9(a) with initial states prepared with an offset. (c) The amplitude precision vs $\bar{n}$ for both SCS (blue circles) and CS (purple diamonds). The dotted lines are linear fit of $\log(\Delta\gamma)=a\log(\overline{n})+b$ with $a$ and $b$ as free parameters. The fitted parameters are $a,b=-0.460(3),-1.311(2)$ for SCS and $a,b=-0.0105(2),-0.480(1)$ for CS.
@@ Appendix E Amplitude estimation using the superposition of coherent states
To illustrate the versatility of our protocol, we demonstrate quantum-enhanced precision for amplitude estimation by simply changing the process unitary in Fig. 2(a) in the main text to a displacement in the imaginary direction $\hat{U}(\gamma)=\exp(i\gamma(\hat{a}^{\dagger}+\hat{a}))$, where $\gamma$ is the displacement amplitude. The process generator is now $\hat{x}=(\hat{a}^{\dagger}+\hat{a})$, making states with high variance $(\Delta\hat{x})^{2}$ useful for amplitude estimation. Here, SCS offer high variance $(\Delta\hat{x})^{2}=2\overline{n}+1$ compared to CS, which have a constant variance $(\Delta\hat{x})^{2}=1$. The illustration is similar to that of phase estimation of Fig. 2(b) in the main text, except the SCS are now displaced in the imaginary direction instead of rotated. SCS with higher $\alpha$ have more rapidly oscillating fringes such that the overlap with the initial state reduces more sharply as $\gamma$ increases.
@@ F.2 Phase estimation
Next, we compare the uncertainty $\Delta\theta$ derived from Bayesian estimation with the Cramér-Rao bound prediction via $\Delta\theta=1/\sqrt{N\,\text{FI}}$ in Fig. 13. The close agreement between the two methods, within the dynamical range, demonstrates the effectiveness of Bayesian inference in achieving near-optimal parameter estimation. The minimal uncertainty $\Delta\theta_{\text{min}}$ as a function of the coherent state amplitude $\alpha$ (or average photon number $\bar{n}$) exhibits a decreasing trend, as shown in Fig. 14, indicating enhanced precision with larger $\bar{n}$. This scaling is consistent with theoretical predictions based on the Cramér-Rao bound.